Given the following sets of information, calculate the requested quantities for each problem.

  1. [H3O+] = 4.1 x 10-4 M. Find pH, and [OH-]
  2. [OH-] = 6.6 x 10-9 M. Find pOH, and [H3O+]
  3. pH = 9.0. Find [H3O+], [OH-] and pOH
  4. pOH = 12.5. Find [OH-], and [H3O+]
  5. pOH = 7.0. Find [H3O+], [OH-] and pH 
1
Expert's answer
2020-11-10T14:01:40-0500

Solution.

"pH = -lg[H3O^+]"

"pOH = -lg[OH^-]"

"pH + pOH = 14"

"pH = 14-pOH"

"pOH = 14-pH"

1.

pH = 3.39

pOH = 10.61

"[OH^-] = 10^{-pOH}"

"[OH^-] = 2.45\times 10^{-11}"

2.

pOH = 8.18

pH = 5.82

"[H3O^+] =10^{-pH}"

"[H3O^+] = 1.51 \times 10^{-6}"

3.

pOH = 5.0

"[H3O^+] = 10^{-9}"

"[OH^-] = 10^{-5}"

4.

pH = 1.5

"[H3O^+] = 3.16 \times 10^{-2}"

"[OH^-] = 3.16 \times 10^{-13}"

5.

pH = 7.0

"[H3O^+] = 10^{-7}"

"[OH^-] = 10^{-7}"

Answer:

1.

pH = 3.39

"[OH^-] = 2.45\times 10^{-11}"

2.

pOH = 8.18

"[H3O^+] = 10^{-9}"

3.

pOH = 5.0

"[H3O^+] = 10^{-9}"

"[OH^-] = 10^{-5}"

4.

"[H3O^+] = 3.16 \times 10^{-2}"

"[OH^-] = 3.16 \times 10^{-13}"

5.

pH = 7.0

"[H3O^+] = 10^{-7}"

"[OH^-] = 10^{-7}"